OpenStudy (anonymous):

Solve for x: −4|x + 5| = −16 X= 21/4,x=-11/4 x=-1,x=9 x=-1,x=-9 no solution please help:)

OpenStudy (anonymous):

start by dividing both sides by \(-4\) to get \[|x+5|=4\]

OpenStudy (anonymous):

then solve two separate equations \(x+5=4\) and also \(x+5=-4\) they each only take one step: subtract \(5\) from both sides

OpenStudy (anonymous):

thank you! @satellite73

OpenStudy (anonymous):

yw